You have deposited $42,000 into an account that will earn an interest rate of 12% compounded semiannually. How much will you have in this account at the end of seven years?

$92,848.62

$71,218.48

$94,957.97

$113,949.56
Answer #1

Hi

Here present value P= $42,000

Interest rate r= 12%/2 = 6%

Time t= 7*2 =14 semi annual

Future value F= P*(1+r)^t

F= 42000*1.06^14

F= $94,957.97

Hence 3rd option is correct here

Thanks